OpenStudy (anonymous):

show steps 5n + 4 = 7(n+1) - 3n

OpenStudy (precal):

5n+4=7n+7-3n distribute 7 5n+4=4n+7 combine like terms n+4=7 move 4n to the other side by subtracting it n=3 move 4 to the other side by subtracting it